OpenStudy (anonymous):

LCD QUESTION!!! What is the LCD of 5/12 and negative 9/16? ________ @phi @iGreen @uChezzy @bibby @perl @PaulaLovesSchool13

OpenStudy (phi):

can you factor 12 ?

OpenStudy (ipurple):

Ok, can you factor the 5/12 and 9/16?

OpenStudy (anonymous):

I'll find a link to help you in the future

OpenStudy (anonymous):

12,24,36,48,60,72,84,96?

OpenStudy (phi):

I was thinking 12 = 2*2*3 now do 16 the same way

OpenStudy (anonymous):

16,32,48,64,80,96,112?

OpenStudy (anonymous):

Psttttt....when you are done and got your answer take a peek here: http://www.helpwithfractions.com/math-homework-helper/least-common-denominator/

OpenStudy (phi):

ok, you are look at multiples. That way works. You pick the first number that is on both lists

OpenStudy (anonymous):

96?

OpenStudy (phi):

96 is on both. but there is a smaller number

OpenStudy (anonymous):

48?

OpenStudy (phi):

yes

OpenStudy (anonymous):

medaled!!

OpenStudy (anonymous):

Megan, dear, check the link out, okay?

OpenStudy (anonymous):

ok

OpenStudy (phi):

fyi, the other way is write 12= 2*2*3 16= 2*2*2*2 12 needs 2 2's and 1 "3" so we start with 2*2*3 16 needs 4 2's. we have 2 2's. add in 2 more: 2*2*3*2*2 that will be the answer also.